OpenStudy (anonymous):

Determine whether the points (1, -2) and (0, 1) lie on the graph of x^2 + y^2 - 3x + y = 0.

OpenStudy (anonymous):

would you like me to show you how to do this?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

replace x by 1 and y by -2 to see if it works

OpenStudy (anonymous):

\[1^2+(-2)^2-3\times 1+(-2)=0?\]

myininaya (myininaya):

say no

OpenStudy (anonymous):

\[1+4-4-2=0\] \[-1=0\] NO

OpenStudy (anonymous):

just say no

myininaya (myininaya):

1+4-3-2 =5-3-2 =2-2 =0 wait say yes

myininaya (myininaya):

lol